Ch I. Keys to the Study of Chemistry 1. Unit conversion. 2.5 pts) 2A. Convert 862.35 torr to atm and mmHg (I atm 760 mmHg) 2B. Convert 365.00 Liters to ml. 0 mL). 2C. Convert 4.536 g to mg and kg (I kg 10 g. g 10 mg). 2D. Convert 25.12 oC to Kelvin. 2E. Convert 25.12 OC to Fahrenheit (in F) 9/5 T (in oC)+ 32) 2F. Convert 1.65 m to nm (1 m 0 nm). 2G. Convert 75 miles to m (1 mile .6 km. 1 km 10 m). 2H. Convert 2.54 m to inch (I in 2.54 cm. 1 m 00 cm) 21. Convert 2.54 gal to liters (IGal 3.80 L) 2J. Convert 2.54 Hz to s
Answer #1

2A)

Given P = 862.35 torr

1 torr = 0.001316 atm

then

862.35 torr = 0.001316 x 862.35 atm = 1.135 atm

1 atm = 760 mmHg

then

1.135 atm = 760 x 1.135 mmHg = 862.49 mmHg

2B)

365.00 L to mL

1 L = 103 mL

then

365 L = 365 x 103 mL = 365000 mL

2C)

4.536 g to mg and Kg

1 g = 103 mg

then

4.536 g = 4.536 x 103 mg = 4536 mg

1 Kg = 103 g

then 1 g = 10-3 Kg

so 4.536 g = 4.536 x 10-3 Kg = 0.0045 Kg

2D)

25.12 0C to Kelvin

We know that 0 0C = 273.15 K

then

25.12 0C = 25.12 + 273.15 = 298.27 0C